OpenStudy (anonymous):

if x^2 ∝ y^2 then is x ∝ y ???if not,,plz xplain!

OpenStudy (amistre64):

a linear equation amounts to the highest degree being a "1": y = x+3 when you start sqaureing things, you start to bend the line.... x^2 and y^2 bend and cross the axises twice. making them conic in nature.

OpenStudy (amistre64):

or so i think....

OpenStudy (anonymous):

i askd smthng else sir!

OpenStudy (amistre64):

im sure you did...... :)

OpenStudy (anonymous):

"If x^2 ∝ y^2 then is x ∝ y?" No, now flutter off.

OpenStudy (anonymous):

Wow. Newton.. How about you take your attitude elsewhere. Just because someone is mistaken doesn't mean you get to be uncivil.

OpenStudy (amistre64):

isnt that a german word??

OpenStudy (anonymous):

When someone makes about 5 threads with the same question, then spams other threads with it also, you think they still deserve respect? Nice.

OpenStudy (anonymous):

I think that there are more civil ways to tell someone that they don't need to repeatedly repost the same thing over and over, just as there are less civil ways I could be explaining this to you. If we don't treat each other with respect things quickly devolve into the kind of "NO U" nonsense you see everywhere else online.

OpenStudy (anonymous):

i m sry guys...m jst 16 nd a bit confused

OpenStudy (anonymous):

I could have explained it several ways, like telling him that an equation in x^2 and y^2 only will have symmetry in both axis, so can, by definition, not be a linear equation. Or another way. .... If he'd asked once. Polpak, you can have your 81 fans, and I'll have my superior Mathematical abilities? Deal?

OpenStudy (anonymous):

sry sir .........sir newton!

OpenStudy (anonymous):

You also have a superior attitude, which is the part I take issue with.

OpenStudy (anonymous):

i'll get it cleared somewhere else by the way....u guyz dont fight plzzzzzzz

OpenStudy (anonymous):

dats why i askd polpak his age?

OpenStudy (anonymous):

thank u 4 ur tym!!!

OpenStudy (anonymous):

i dont know hyperbola!

OpenStudy (anonymous):

so cant get ur terms of usage in xplanation....was a bit complex frm ur syd!!!!

OpenStudy (anonymous):

i m an indian and we give due respect to ol even if they dont give us!!!!

OpenStudy (anonymous):

Sketch the graphs.

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

x^2 = y^2, therefore x = ±y , so this graph will consist of 2 lines, not one (x=y in the first quadrant, reflected into all the others, which you could see as 2 lines, x = y and x = -y)

OpenStudy (amistre64):

..... i think that means I was right lol ;)

OpenStudy (anonymous):

The graphs only start curving into hyperbola if y^2 = mx^2 + k, not if y^2 = mx^2 i.e. y^2 propto x^2

OpenStudy (amistre64):

true....